OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. (7 points) a16 = 21, a17 = -1 (a)an = 351 + 22(n + 1) (b)an = 351 - 22(n - 1) (c)an = 351 + 22(n - 1) (d)an = 351 - 22(n + 1)

OpenStudy (anonymous):

Do you know the general equation for the nth term of a sequence?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a _{n}=a _{1}+d(n-1)\] a1 = the first term d = common difference (the space between consecutive terms) n = term number

OpenStudy (anonymous):

You have terms 16 and 17, so you can subtract them to get d. \[d=a _{17}-a _{16}\]

OpenStudy (anonymous):

21-(-1)=22?

OpenStudy (anonymous):

other way around. the 17th term is -1 and the 16th is 21

OpenStudy (anonymous):

-1-21=(-22)?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So it's B!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (anonymous):

yw